OpenStudy (anonymous):

f(x)=(x^2-4)/(x^2+1) Find any asymptotes, extrema, intercepts, and points of inflection.

OpenStudy (anonymous):

I've found an asymptote at x=1, x-intercepts at -2 and 2, and a y-intercept at -4. I'm having trouble finding the extrema...

OpenStudy (anonymous):

\[f'(x)= (x^2+1)(2x)-(x^2-4)(2x)/(x^2+1)^2\]

OpenStudy (anonymous):

which then equals \[2x^3+2x-2x^3+8x/(x^2+1)^2\]

OpenStudy (anonymous):

\[\frac{x^2-4}{x^2+1}=1-\frac{5}{x^2+1} \]

OpenStudy (anonymous):

What is that?

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

\[\lim_{x\to \infty } \, \left(1-\frac{5}{x^2+1}\right)=1 \] Same limit as x approaches -Infinity

OpenStudy (anonymous):

Set the 2nd derivative of f(x) to zero and solve for x. Plug the x's into f(x) to find the corressponding y's for the inflection points.\[\left\{\pm \frac{1}{\sqrt{3}},-\frac{11}{4}\right\} \]

OpenStudy (anonymous):

Okay, thank you! But I am having trouble finding the extrema still...

OpenStudy (anonymous):

It appears that I have a minimum at x=-4, but I need to find it by doing the math.

OpenStudy (anonymous):

Set the first derivative to zero\[\frac{10 x}{\left(x^2+1\right)^2}=0 \]and solve for x.\[x=0 \] if x=0 then the corresponding y is -4 using\[\frac{x^2-4}{x^2+1} \]

OpenStudy (anonymous):

THANK YOU! I don't know why I couldn't do that earlie!

OpenStudy (anonymous):

You are very welcome.

OpenStudy (anonymous):

Is \[-2x(x^2+2)/(x^2+1)^2\] the correct second derivative?

OpenStudy (anonymous):

The second derivative is:\[-\frac{10 \left(3 x^2-1\right)}{\left(x^2+1\right)^3} \]